# R-Bot: An LLM-based Query Rewrite System

### v1

@@ 4.1. Rewrite Rule Specification Preparation
Figure 3. Transforming Rule Code into Rule Specification.
@@ 4.1.1. Transforming Rule Code into Rule Specification
Given that some query rewrite engines, such as Apache Calcite (cal, 2024), are not accompanied by comprehensive documentation, we are compelled to decipher the rewrite rules directly from the raw code. This process involves navigating through complex code structures that include intricate nested calls. To address this challenge, we introduce a hierarchical rule code summarization method, as illustrated in Figure 3. Our approach begins with the construction of a rule code structure tree, emanating from the rule’s main function. In this tree, each node represents a symbol declaration (e.g., functions, variables, classes), while each edge denotes a symbol reference relationship. Progressing through the structure, we methodically summarize the declaration code, moving from simple to more complex elements and clarifying symbol references using previously summarized symbols. With the comprehensive summary of the root node at our disposal, we guide LLM to convert this information into a standardized rewrite rule specification.
Rule Code Structure Analysis. Given the symbol references of the rule code, it is crucial to clarify the symbol declarations before summarizing the rule code. We first build a root node representing the main function of the rule. Then, we use code analysis tools (e.g., JavaSymbolSolver (jav, 2018)) to associate the symbols in the main function with their corresponding declarations, which are children nodes of the root node. If the declaration of some nodes also accesses other unseen symbols, we further resolve its symbol references. We recursively expand the nodes until reaching built-in symbols. In this way, we obtain a rule code structure tree, where each node represents a symbol declaration and each edge represents a symbol reference relationship.
Hierarchical Rule Code Summarization. The complex rule code structure poses two challenges for LLM code summarization. First, given the necessity of every declaration in the rule code structure, if we directly input them to LLM, the long context can greatly degrade LLM performance (Liu et al., 2024a; Li et al., 2023). Second, the considerable width and depth of the structure (e.g., tens) further increase the reasoning complexity for LLM. We thus propose a hierarchical rule code summarization method. First, if a node declaration already has detailed comments, there is no need to summarize the node. Second, for the leaf node without comments, it can be directly summarized by LLM due to the simple code structure and absence of unfamiliar symbols. Third, for the non-leaf node whose children are already summarized, the symbol references in the declaration code can be clarified by its children summaries. Specifically, we insert the symbol summaries as comments into the declaration code, which enables LLM to accurately summarize the code. We repeat the process until we obtain the summary of the root, which represents the whole rule code.
@@ 4.1.2. Transforming Rewrite Document into Rule Specification
Rule Extraction. We use LLM to extract rewrite rules from the rewrite documents in two steps. First, if we directly input all the documents to LLM, it often overlooks important details in the middle of the extremely long context (e.g., 100k) (Liu et al., 2024a; Li et al., 2023). We thus split the documents into structured blocks (e.g., sections, sub-sections) that each can be effectively processed by LLM. Second, we instruct LLM to extract rewrite rules from the split blocks. To mitigate the hallucination problem (Rawte et al., 2023), we require LLM to locate supporting content in the source document. If no such content can be found, we can deem the extraction low quality and repeat LLM extraction.
Rule Clustering. To identify pertinent rules among different chapters and documents (e.g., condition push-down involved in both PostgreSQL and MySQL documents), we adopt a clustering algorithm based on Gaussian Mixture Models (GMMs) (Sarthi et al., 2024). Specifically, the semantics of each rule is represented as a embedding vector (e.g., using multi-qa-mpnet-base-cos-v1 (Reimers and Gurevych, 2019; mul, 2024)). It then fits the vector distribution with a weighted combination of $M$ Gaussian distributions. Each distribution forms a cluster, assigning rules to the cluster with the highest density of their embeddings. To address the limitations of GMMs in high-dimensional spaces, it adopts Uniform Manifold Approximation and Projection, which reduces dimensionality by approximating the local manifold of the data (McInnes et al., 2018). Besides, to determine the optimal cluster number $M$, it adopts Bayesian Information Criterion, which aims to maximize the overall distribution likelihood $\hat{L}$ while minimizing the cluster number (Schwarz, 1978). This is achieved by minimizing $ln(N)m-2ln(\hat{L})$, where $N$ is the number of rules, and $m$ is proportional to the cluster number $M$.
Rule Specification Regularization. We use LLM to obtain standardized rule specifications from the rule clusters. First, we summarize each cluster using LLM with a prompt $p_{summ}=$“Given the rewrite rule components, your task is to summarize them into one paragraph, and your summary should include as many details as possible.” Then, we further use LLM to transform the summary to a regularized rewrite rule specification, by extracting the conditions and transformations of the rule.
@@ 4.2. Rewrite Q&A Preparation
The rewrite Q&A showcases query rewrite of how to transform a particular SQL query to an optimized one, which is composed of two parts: $(i)$ “question” that denotes a query rewrite request; $(ii)$ “answer” that provides the query transformations for the question. There are plenty of rewrite Q&As within online database community forums (e.g., millions at Stack Overflow (sta, 2024)), which may cover more flexible and practical cases beyond the rewrite rules.
Since the website Q&As are mixed of diverse topics and qualities, we use a hybrid method to filter out the high-quality Q&As on query rewrite. First, we select Q&As by their question tags (e.g., “query-optimization”) and community comments (e.g., a Stack Overflow score higher than 3). Second, we use LLM to examine the topics of these Q&As, and select those related to query rewrite.
5. Hybrid Structure-Semantics Evidence Retrieval
@@ 5.2.1. Structure-Aware SQL Query Embedding
Existing query embedding models (Tang et al., 2022; Zhao et al., 2022, 2023; ope, 2024) cannot be directly applied for embedding the query structure for query rewrite due to limitations in two key aspects. On one hand, the query-specific embeddings mostly adopt small-scale neural networks (e.g., lower than 0.1B) fine-tuned on a specific dataset (Zhao et al., 2022, 2023) or a specific workload (Tang et al., 2022), and thus struggle to generalize to arbitrary scenarios. On the other hand, the general text embeddings (e.g., text-embedding-3-small (ope, 2024)) cannot be directly used for query structure embedding, which are sensitive to irrelevant query information for three reasons. $(i)$ The identifiers (e.g., table/column names) and literals in the query often disrupt characterization of the query structure. For example, while renaming the schema in a query can maintain the integrity of the query rewrites, this action results in a version with significantly altered semantics and it generates an embedding distinctly different from that of the original query; $(ii)$ Since query rewrites typically focus on transforming only certain parts of the query (e.g., a sub-query), disregarding other irrelevant parts can help clarify the query structure; $(iii)$ The text embedding models cannot guarantee commutative invariance, implying they may generate dissimilar embeddings for expressions such as “$p_{1}\ and\ p_{2}$” and “$p_{2}\ and\ p_{1}$”. These can be addressed through carefully designed textual transformations.
@@
Dataflow Based Query Template Embedding. To effectively capture the essential SQL features for query rewrite, we initially reformulate the SQL query into composite dataflows that detail operations on the involved identifiers. Specifically, given the structure and semantics of a query $q$, a dataflow $\delta$ is a list of SQL operations sequentially performed on one or multiple tables or columns in the query, which corresponds to an SQL segment of the query, e.g., logical expression, mathematical expression, “WHERE” clause, “GROUP BY” clause, “FROM” clause, “JOIN” clause, “SELECT” list, sub-query, etc. For instance, we show sample dataflows for the input query in Figure 1, concentrating on the column “$bonus.sal$”:
$\{sal,SELECT\ MAX(sal),emp.sal=ANY\ (SELECT\ MAX(sal)\ ...)\}$
where “$sal$” is first input from “$bonus$” table, projected with an aggregate function “$MAX(\cdot)$”, and then compared with the column “$emp.sal$” from the outer query.
@@
To ensure that the generated query template remains unaffected by schema renaming, we propose an isolated masking method. If a query template involves multiple identifiers, swapping any two identifiers alters its semantics and disrupts the query template embedding. Thus, we derive multiple query templates from the query, with each template preserving only one identifier while masking the others. We then refine the query templates in three steps. First, since a complex query containing many identifiers can derive a large number of query templates, we opt to select only a small part of them as representative. Specifically, if the SQL operations associated with an identifier are potential to be matched by some rewrite rules, we select the corresponding query template as representative, since it concentrates on those potential SQL operations on the identifier. Second, we simplify the query template by eliminating the SQL operations that do not involve the key identifier. Third, for the commutative SQL operands, we sort them lexically, thereby rendering the query template resistant to potential commutations.
Step 1: Potential Identifier Selection. Considering the SQL operations associated with the identifier, we observe that if the identifier appears multiple times in some mutually exclusive dataflows, it can possibly trigger query rewrite by utilizing the correlations of the dataflow operations. For instance, the twice appearance of column “$a$” in expressions “$a<b$” and “$a=5$” can indicate constant folding, which replaces “$a<b$” with “$5<b$”. Then, given an SQL query, we propose two ways to count the frequency of identifier occurrences. $(i)$ Column appearances: By traversing the query, we count how often each column appears. Note that columns appearing as direct projections (e.g., “SELECT a”) are not counted, as this does not involve any meaningful SQL operations. $(ii)$ Table appearances: We count the appearances of the table where the table is utilized either in the “FROM” clause or within a “JOIN” operation. With the frequency of tables and columns, we focus on identifiers that appear more than once within the query. We then select their corresponding query templates as representative, ensuring we capture the most significant patterns for query rewrite.
@@
One-Hot Embedding for Matched Rule Specifications. Given an SQL query, its matched rule specifications also reflect the query structure. For instance, queries matching “SUB_QUERY_TO_JOIN” rule likely contain similar query structures surrounding particular sub-queries. As shown in Section 5.1, we first examine the rule specifications with their associated matching functions. Following this, we construct a one-hot encoding from the rule matching results, where each position corresponds to a rule specification. If a rule specification matches the query, its corresponding position is ‘1’; ‘0’ otherwise. The one-hot embedding for matched rule specifications can capture local structural information indicated by the rule condition, which is complementary to the global structures captured by query template embedding.
@@ 5.2.2. Structure-Semantics Q&A Retrieval
First, we build a hybrid structure-semantics embedding index for Q&A repository in three steps. $(i)$ We extract the queries in the Q&As, generate query templates, and embed them with pre-trained text embedding (e.g., text-embedding-3-small (ope, 2024)). We also match the queries with rule specifications to build one-hot embedding for matched rule specifications. $(ii)$ We refine the Q&A by preserving the semantics of query rewrite and eliminating irrelevant text with the help of LLM, then embedding the streamlined information using a pre-trained text embedding model (e.g., text-embedding-3-small (ope, 2024)). $(iii)$ We normalize the three embeddings as unit vectors, concatenate them to form a holistic embedding, and insert them into the index. Next, for an online SQL query, we similarly generate its embedding and identity the most relevant Q&As with top-$k$ embedding similarities.
@@ 6. Step-by-Step LLM Rewrite
$(i)$ Since the rules have matching functions, we first filter the rules whose matching functions are satisfied by the input query. However, the matched rules exhibit varied relevance to the query with regards to their actual transformation. Specifically, the rules of the query rewrite engine can be classified into two types of transformations: the normalization rule (e.g., “SUB_QUERY_TO_JOIN”), which nearly always reduces query cost; and the exploration rule (e.g., “AGGREGATE_JOIN_TRANSPOSE”), which transforms the query but does not consistently result in cost reduction (Taft et al., 2020). We classify the rules based on expert experience, and assign relevance scores to the matched rules based on their transformation types. Initially, every rule has a score of $\infty$- ∞. Then, the matched normalization rules are assigned +$\infty$+ ∞ as closely relevant, and the matched exploration rules are assigned 00 as weakly relevant.
@@
Step 2: Recipe-based Rule Selection and Ordering. Derived recipes offer more detailed and comprehensive guidance for rewriting the input query when compared to the basic rule specifications and Q&As. Therefore, we refine the initially arranged rule sequence to align it more closely with these recipes. This refinement process is divided into three sub-steps.
$(i)$ We first refine rule selection with recipes. A straightforward method is to assess the relevance between each rule and each recipe, excluding those rules deemed irrelevant. However, this method faces two significant drawbacks. Firstly, it overlooks rule dependencies, consequently excluding indirectly utilized rules that become relevant only after another rule is applied. Secondly, this method incurs high costs, especially since it lies on the critical path for online query rewrite. To address these limitations, we propose a batch filtering method. Specifically, we first feed the rewrite recipe and a batch of rules to LLM, and require LLM to select the most appropriate rules aligned with the rewrite recipe, using the prompt $p_{select\_rule}=$“Given the input query and the rewrite rules, you should evaluate whether the rules can be applied to rewrite the query in the context of the recipe.” Next, we feed the previously selected rules along with the next batch of rules into LLM for further selection. We continue this process iteratively until all rules have been evaluated. In this way, we not only accelerate the LLM selection process, but also consider the rule dependency within the selected rules.
@@
LearnedRewrite
67.12
($\downarrow$36.0%)
10.60
($\downarrow$0.0%)
300.00
($\downarrow$0.0%)
33.67
($\downarrow$10.8%)
4.67
($\downarrow$11.6%)
57.60
($\downarrow$80.8%)
12.63
($\downarrow$88.5%)
5.24
($\downarrow$90.7%)
48.47
($\downarrow$83.8%)
GPT-3.5
85.98
($\downarrow$18.0%)
10.60
($\downarrow$0.0%)
300.00
($\downarrow$0.0%)
37.75
($\downarrow$0.0%)
5.36
($\uparrow$1.5%)
300.00
($\downarrow$0.0%)
55.41
($\downarrow$49.5%)
22.74
($\downarrow$59.6%)
230.99
($\downarrow$23.0%)
GPT-4
67.10
($\downarrow$36.0%)
10.60
($\downarrow$0.0%)
300.00
($\downarrow$0.0%)
37.77
($\uparrow$0.0%)
4.92
($\downarrow$6.8%)
300.00
($\downarrow$0.0%)
60.86
($\downarrow$44.5%)
20.06
($\downarrow$64.4%)
300.00
($\downarrow$0.0%)
R-Bot (GPT-3.5)
56.24
($\downarrow$46.4%)
10.60
($\uparrow$0.0%)
300.00
($\downarrow$0.0%)
26.15
($\downarrow$30.7%)
4.61
($\downarrow$12.7%)
35.25
($\downarrow$88.2%)
37.71
($\downarrow$65.6%)
8.37
($\downarrow$85.1%)
65.67
($\downarrow$78.1%)
R-Bot (GPT-4)
57.60
($\downarrow$45.1%)
10.37
($\downarrow$2.2%)
300.00
($\downarrow$0.0%)
25.35
($\downarrow$32.9%)
4.58
($\downarrow$13.2%)
17.17
($\downarrow$94.3%)
12.45
($\downarrow$88.6%)
5.04
($\downarrow$91.0%)
48.30
($\downarrow$83.9%)

### v2

@@ 4.1.1. Transforming Rule Code into Rule Specification
Given that some query rewrite engines, such as Apache Calcite (cal, 2024), are not accompanied by comprehensive documentation, we are compelled to decipher the rewrite rules directly from the raw code. This process involves navigating through complex code structures that include intricate nested calls. To address this challenge, we introduce a hierarchical rule code summarization method, as illustrated in Figure 3. Our approach begins with the construction of a rule code structure tree, emanating from the rule’s main function. In this tree, each node represents a symbol declaration (e.g., functions, variables, classes), while each edge denotes a symbol reference relationship. Progressing through the structure, we methodically summarize the declaration code, moving from simple to more complex elements and clarifying symbol references using previously summarized symbols. With the summary of the root node at our disposal, we guide LLM to convert this into a standardized rewrite rule specification.
Rule Code Structure Analysis. Given the symbol references of the rule code, it is crucial to clarify the symbol declarations before summarizing the rule code. We first build a root node representing the main function. Then, we use code analysis tools (e.g., JavaSymbolSolver (jav, 2018)) to associate the symbols with their corresponding declarations, which are children nodes of the root node. If the declaration of some nodes also accesses other unseen symbols, we further resolve its symbol references. We recursively expand the nodes until reaching built-in symbols. In this way, we obtain a rule code structure tree, where each node represents a symbol declaration and each edge represents a symbol reference relationship.
Hierarchical Rule Code Summarization. The complex rule code structure poses two challenges for LLM summarization. First, due to the relevance of nearly every declaration in the code, if we directly input them to LLM, the long context can greatly degrade LLM performance (Liu et al., 2024a; Li et al., 2023). Second, the substantial width and depth of the code structure (e.g., tens of nodes) further increase the reasoning burden. To address these issues, we propose a hierarchical rule code summarization method. First, if a node declaration already has detailed comments, summarization is unnecessary. Second, leaf node without comments can be directly summarized by LLM due to the simple code and absence of unfamiliar symbols. Third, for the non-leaf node whose children are already summarized, the symbol references in the declaration code can be clarified by its children summaries. Specifically, we insert the symbol summaries as comments into the declaration code, which enables accurate LLM summarization. This process is repeated recursively until the root node is summarized, yielding a summary of the entire rule code.
@@ 4.1.2. Transforming Rewrite Document into Rule Specification
Figure 3. Transforming Rule Code into Rule Specification.
@@
Rule Extraction. We use LLM to extract rewrite rules from the rewrite documents in two steps. First, if we directly input all the documents to LLM, it often overlooks important details in the middle of the extremely long context (e.g., 100k) (Liu et al., 2024a; Li et al., 2023). We thus split the documents into structured blocks (e.g., sections, sub-sections) that each can be effectively processed by LLM. Second, we instruct LLM to extract rewrite rules from the split blocks. To mitigate the hallucination problem (Huang et al., 2023), we require LLM to locate supporting content in the source document. If no such content can be found, we can deem the extraction low quality and repeat LLM extraction.
Rule Clustering. To identify pertinent rules among documents (e.g., condition push-down involved in both PostgreSQL and MySQL documents), we cluster the rules based on rule semantics. Specifically, we embed each rule into vectors (e.g., using multi-qa-mpnet-base-cos-v1 (Reimers and Gurevych, 2019; mul, 2024)), and cluster them using Gaussian Mixture Models (Sarthi et al., 2024). To enhance clustering in high-dimensional space, we apply Uniform Manifold Approximation and Projection for dimensionality reduction by approximating local data manifold (McInnes et al., 2018).
Rule Specification Regularization. We use LLM to obtain standardized rule specifications from rule clusters. First, we summarize each cluster using LLM with prompt $p_{summ}=$“Given rewrite rule components, your task is to summarize them into one paragraph, and your summary should include as many details as possible.” Then, we use LLM to transform the summary to regularized rewrite rule specification, by extracting rule condition and transformation.
@@ 4.2. Rewrite Q&A Preparation
The rewrite Q&A showcases how to rewrite a particular SQL query to an optimized one, which is composed of two parts: $(i)$ “question” that denotes a query rewrite request; $(ii)$ “answer” that provides the query transformations for the question. There are plenty of rewrite Q&As within online database community forums (e.g., millions at Stack Overflow (sta, 2024)), often covering practical cases beyond standard rewrite rules. To filter high-quality rewrite Q&As from these mixed sources, we use a hybrid method: first selecting by question tags (e.g., “query-optimization”) and community feedback (e.g., Stack Overflow score higher than 3), and then using LLM to verify their relevance to query rewrite.
5. Structure-Semantics Retrieval
@@ 5.2.1. Structure-Aware SQL Query Embedding
Existing query embedding models (Tang et al., 2022; Zhao et al., 2022, 2023; ope, 2024) cannot be directly applied for embedding query structure for query rewrite due to two key limitations. On one hand, the query-specific embeddings mostly adopt small-scale neural networks (e.g., lower than 0.1B) fine-tuned on narrow datasets (Zhao et al., 2022, 2023) or workloads (Tang et al., 2022), limiting their generalizability. On the other hand, the general text embeddings (e.g., text-embedding-3-small (ope, 2024)) fail to effectively represent query structure, as they are sensitive to irrelevant information: $(i)$ The identifiers (e.g., table/column names) and literals in the query often disrupt characterization of query structure. For example, while renaming the schema in a query can maintain the integrity of the query rewrites, this action results in a version with significantly altered semantics and it generates an embedding distinctly different from that of the original query; $(ii)$ Since query rewrites typically focus on transforming only certain parts of the query (e.g., a sub-query), disregarding other irrelevant parts can help clarify the query structure; $(iii)$ The text embedding models cannot guarantee commutative invariance, implying they may generate dissimilar embeddings for expressions such as “$p_{1}\ AND\ p_{2}$” and “$p_{2}\ AND\ p_{1}$”. These can be addressed through carefully designed textual transformations.
@@
Dataflow Based Query Template Embedding. To capture the essential SQL features for query rewrite, we initially reformulate the SQL query into composite dataflows that detail operations on the involved identifiers. Specifically, given a query $q$, a dataflow $\delta$ is a list of SQL operations sequentially performed on one or multiple tables or columns in the query, which corresponds to an SQL segment of the query, e.g., logical expression, mathematical expression, “WHERE” clause, “GROUP BY” clause, “FROM” clause, “JOIN” clause, “SELECT” list, sub-query, etc. For instance, we show sample dataflows for the input query in Figure 1, concentrating on the column “$bonus.sal$”: $\{sal,SELECT\ MAX(sal),emp.sal=ANY\ (SELECT\ MAX(sal)\ ...)\}$ where “$sal$” is first input from “$bonus$” table, projected with an aggregate function “$MAX(\cdot)$”, and then compared with the column “$emp.sal$” from the outer query.
@@
To ensure query templates are robust to schema renaming, we propose an isolated masking method. If a query template involves multiple identifiers, swapping any two identifiers alters its semantics and disrupts the embedding. Thus, we derive multiple query templates from the query, with each template preserving only one identifier while masking the others. We then refine the templates in three steps. First, to limit the number of templates derived from complex queries, we retain only those where the SQL operations associated with the identifier are likely to match certain rewrite rules. Second, we simplify the query template by eliminating SQL operations that do not involve the key identifier. Third, for the commutative SQL operands, we sort them lexically, thereby rendering the template invariant to operand order.
Step 1: Potential Identifier Selection. Considering the SQL operations associated with the identifier, we observe that multiple appearances of the same identifier in mutually exclusive dataflows may prompt query rewrites by leveraging operation correlations. For instance, the twice appearance of column “$a$” in “$a<b$” and “$a=5$” can indicate constant folding, replacing “$a<b$” with “$5<b$”. Then, given an SQL query, we propose two ways to measure identifier frequency: $(i)$ Column appearances: By traversing the query, we count how often each column appears. Note that columns appearing as direct projections (e.g., “SELECT a”) are not counted, as this does not involve any meaningful SQL operations. $(ii)$ Table appearances: We count the appearances of the table which is utilized either in the “FROM” clause or within a “JOIN” operation. With the frequency of tables and columns, we focus on identifiers that appear more than once within the query. We then select their corresponding query templates as representative, ensuring we capture the most significant patterns for query rewrite.
@@
One-Hot Embedding for Matched Rule Specifications. Given an SQL query, its matched rule specifications also reflect query structure. For instance, queries matching “SUB_QUERY_TO_JOIN” rule likely exhibit similar structures around sub-queries. As shown in Section 5.1, we first examine the rule specifications with their associated matching functions. Following this, we construct a one-hot encoding from the rule matching results, where each position corresponds to a rule specification. If a rule specification matches the query, its corresponding position is ‘1’; ‘0’ otherwise. The one-hot embedding for matched rule specifications can capture local structural information indicated by the rule condition, complementary to the global structures captured by query template embedding.
@@ 5.2.2. Structure-Semantics Q&A Retrieval
First, we build a structure-semantics embedding index for Q&A repository in three steps. $(i)$ We extract the queries in the Q&As, generate query templates, and embed them with pre-trained text embedding (e.g., text-embedding-3-small (ope, 2024)). We also match the queries with rule specifications to build one-hot embedding for matched rule specifications. $(ii)$ We refine the Q&A by preserving the semantics of query rewrite and eliminating irrelevant text with the help of LLM, then embedding the streamlined information using a pre-trained text embedding model (e.g., text-embedding-3-small (ope, 2024)). $(iii)$ We normalize the three embeddings as unit vectors, concatenate them to form a holistic embedding, and insert them into the index. Next, for an online SQL query, we similarly generate its embedding and identify the most relevant Q&As with top-$k$ embedding similarities.
@@ 6. Step-by-Step LLM Rewrite
$(i)$ We first filter rules whose matching functions are satisfied by the input query. However, the matched rules vary in relevance based on their actual transformation. Specifically, the rules of the query rewrite engine can be classified into two types of transformations: the normalization rule (e.g., “SUB_QUERY_TO_JOIN”), which nearly always reduces query cost; and the exploration rule (e.g., “AGGREGATE_JOIN_TRANSPOSE”), which transforms the query but does not consistently result in cost reduction (Taft et al., 2020). We classify the rules based on expert experience, and assign relevance scores to the matched rules based on their transformation types. Initially, every rule has a score of $\infty$- ∞. Then, the matched normalization rules are assigned +$\infty$+ ∞ as closely relevant, and the matched exploration rules are assigned 00 as weakly relevant.
@@
LearnedRewrite 69.60 ($\downarrow$33.6%) 12.26 ($\uparrow$15.7%) 300.00 ($\downarrow$0.0%) 30.47 ($\downarrow$19.3%) 5.28 ($\downarrow$0.0%) 55.02 ($\downarrow$81.7%) 79.07 ($\downarrow$27.9%) 5.24 ($\downarrow$90.7%) 300.00 ($\downarrow$0.0%)
GPT-3.5 85.98 ($\downarrow$18.0%) 10.60 ($\downarrow$0.0%) 300.00 ($\downarrow$0.0%) 37.75 ($\downarrow$0.0%) 5.36 ($\uparrow$1.5%) 300.00 ($\downarrow$0.0%) 55.41 ($\downarrow$49.5%) 22.74 ($\downarrow$59.6%) 230.99 ($\downarrow$23.0%)
GPT-4 67.10 ($\downarrow$36.0%) 10.60 ($\downarrow$0.0%) 300.00 ($\downarrow$0.0%) 37.77 ($\uparrow$0.0%) 4.92 ($\downarrow$6.8%) 300.00 ($\downarrow$0.0%) 60.86 ($\downarrow$44.5%) 20.06 ($\downarrow$64.4%) 300.00 ($\downarrow$0.0%)
R-Bot (GPT-3.5) 55.71 ($\downarrow$46.9%) 10.41 ($\downarrow$1.8%) 300.00 ($\downarrow$0.0%) 26.19 ($\downarrow$30.6%) 4.61 ($\downarrow$12.7%) 35.25 ($\downarrow$88.2%) 37.71 ($\downarrow$65.6%) 8.37 ($\downarrow$85.1%) 65.67 ($\downarrow$78.1%)
R-Bot (GPT-4) 57.60 ($\downarrow$45.1%) 10.37 ($\downarrow$2.2%) 300.00 ($\downarrow$0.0%) 25.35 ($\downarrow$32.9%) 4.58 ($\downarrow$13.2%) 17.17 ($\downarrow$94.3%) 12.45 ($\downarrow$88.6%) 5.04 ($\downarrow$91.0%) 48.30 ($\downarrow$83.9%)
Step 2: Recipe-based Rule Selection and Ordering. Derived recipes offer more detailed guidance for rewriting the input query compared to the basic rule specifications and Q&As. Therefore, we refine the initially arranged rule sequence to align it more closely with these recipes. This process is divided into three sub-steps.
$(i)$ We first refine rule selection with recipes. A straightforward method is to assess the relevance between each rule and each recipe, excluding those rules deemed irrelevant. However, this method faces two significant drawbacks. Firstly, it overlooks rule dependencies, consequently excluding indirectly utilized rules that become relevant only after another rule is applied. Secondly, this method incurs high costs, especially since it lies on the critical path for online query rewrite. To address these limitations, we propose a batch filtering method. Specifically, we first feed the rewrite recipe and a batch of rules to LLM, and require LLM to select the most appropriate rules aligned with the rewrite recipe, using the prompt $p_{select\_rule}=$“Given the input query and rewrite rules, you should evaluate whether the rules can be applied to rewrite the query in the context of the recipe.” Next, we feed the previously selected rules along with the next batch of rules into LLM for further selection. We continue this process iteratively until all rules have been evaluated. In this way, we not only accelerate the LLM selection process, but also consider the rule dependency within the selected rules.